# Semiclassical Simulation of Homogeneous Emitter Ensembles with Local Dissipation

### v1

@@ Abstract
Emitter ensembles constitute a fundamental component in quantum optical technologies, yet efficient and accurate simulation of large ensembles remains challenging. Here, we formulate a truncated Wigner approximation (TWA) for permutation-invariant emitter ensembles subject to local dissipation by sampling stochastic trajectories in an extended phase space encompassing the Bloch sphere. Benchmarks show that the TWA accurately captures dynamics, including nonclassical signatures, with the approximation improving with ensemble size. We demonstrate large-scale simulations of hundreds of interacting ensembles within the TWA to reveal emergent spatial coherence and selective directionality of cooperative emission in a pumped 1D chain, highlighting a practical path to studying extended light-matter systems. Our results expand the scope of scalable simulations of quantum emitter ensembles, establishing a bridge between microscopic models and emergent behavior.
@@
Homogeneous ensembles of indistinguishable emitters play a fundamental role in understanding and engineering open quantum systems. In state-of-the-art experiments featuring atoms [7, 5, 22, 18] and solid-state [60, 86, 38, 53, 27] emitters, collective interactions commonly coexist with local dissipation. While this dissipation must often be mitigated, the interplay with collective dynamics can enable novel dissipative phases of matter [78, 31, 36, 82, 37] and distinct device functionalities enhanced by cooperativity [41, 54, 48, 71]. As experiments enter the mesoscopic regime, exact simulations [10, 62, 58, 81, 30, 78, 20] of minimal, permutation-invariant models are typically constrained to modest system sizes. This limitation motivates more scalable approximations [79] such as cumulant expansions [35, 51], tensor networks [75, 88], and (discrete) phase-space methods [59, 43, 65, 25, 23, 42]. However, obtaining quantitatively controlled predictions at larger system sizes remains challenging, calling for targeted approaches [34, 4].
@@
Semiclassical phase-space methods [45, 57] provide a powerful route to approximate many-body quantum dynamics, notably through the truncated Wigner approximation (TWA) [77, 52]. In the TWA, quantum evolution is represented by an ensemble average of semiclassical trajectories. Initial quantum fluctuations enter through random sampling, and in open systems dissipation contributes stochastic evolution. In the semiclassical limit, the approximation improves systematically, with quantum corrections entering perturbatively [52]. While well-established for bosons [67, 17, 76, 6] and for spin ensembles with purely collective dynamics [21, 74, 24, 1, 33], an analogous semiclassical description of local dissipation in permutation-invariant ensembles has remained limited.
@@
Fig. 1: (Bottom) An ensemble of emitters uniformly couples to a common field. (Top left) The ensemble emits collectively with rate $\Gamma$ and undergoes independent local dissipation with rate $\gamma$. (Top right) In semiclassical simulations, the ensemble spin (large arrow) is sampled by classical phase-space variables (dots). Stochastic trajectories evolve along the Bloch sphere surface under collective processes, while local dissipation drives evolution into the sphere’s interior.
@@
Here, we develop a semiclassical framework for the dynamics of permutation-invariant emitter ensembles (Fig. 1). Building on a quantum-to-classical mapping for variable-length spins [32], we formulate a truncated Wigner approximation in which the entire ensemble is described using just four phase-space variables that extend the usual Bloch sphere description. The additional variables capture changes in spin length induced by local dissipation, which enters through a few stochastic processes acting on semiclassical trajectories. In bad-cavity superradiance, and boundary time crystals, we find the method is accurate for small ensembles of tens of emitters—capturing squeezing and subradiance—and numerically approaches the exact solution for increasing ensemble size. The cost of the TWA scales only linearly with the number of distinct ensembles, enabling simulations corresponding to tens of thousands of emitters interacting through a resonator chain. In this extended nonlinear medium, we illustrate spatially structured dynamics under uniform pumping and a crossover to strong, selectively directional emission in response to a weak probe, highlighting a broad range of emergent cooperative phenomena in light-matter systems accessible with the TWA.
@@
We consider a homogeneous ensemble formed from $N$ two-level emitters indexed by $n$. Coupling to common modes induces coherent evolution under a Hamiltonian $\hat{H}$ and collective dissipation via jumps $\hat{L}^{q}$ built from collective spin operators $\hat{L}^{0}=\hat{J}^{z}=\frac{1}{2}\sum_{n}\hat{\sigma}_{n}^{z},\ \hat{L}^{\pm 1}=\hat{J}^{\pm}=\sum_{n}\hat{\sigma}_{n}^{\pm}$ where $\hat{\sigma}^{0}_{n}\equiv\hat{\sigma}^{z}_{n}$, $\hat{\sigma}^{\pm 1}_{n}\equiv\hat{\sigma}^{\pm}_{n}$ are the Pauli-$z$ and ladder operators of emitter $n$. Including statistically identical, independent dissipation, the system evolves under the Born-Markov master equation ($\hbar=1$),
@@
To perform a semiclassical expansion of Eq. (1), we exploit invariance under permutation of emitters. Here, the ensemble dynamics can be equivalently expressed in the angular-momentum representation [62, 3, 81], where any permutation-invariant density matrix takes the form,
@@
In this representation, $|JM\rangle$ denotes collective kets labeled by eigenvalues $J(J+1)$ and $M$ of total spin squared $\hat{J}^{2}$ and inversion $\hat{J}^{z}$ respectively, and elements $|JM\rangle\langle JM^{\prime}|$ are identified with the corresponding Dicke-state outer products in the full Hilbert space, averaged over permutation degeneracy [61]. The coefficients $\rho_{JMM^{\prime}}$ then completely specify the state, with $\hat{\rho}$ block diagonal in $J$ and collective operators acting purely as rotations in each block. Local decoherence induces population transfer between distinct sectors $J\leq N/2$ and, crucially, is expressed in the angular-momentum representation using only three effective jump operators [34] for each $q$:
@@
Acting on elements in Eq. (2), the jump operators $\hat{l}^{jq}$ [61], for j=1,0,1j=-1,0,1, couple both inversion $M\to M+q$ and total spin number $J\to J+j$ to neighboring sectors.
We establish the quantum–classical correspondence (details in [61]) for operators in Eq. (1) by applying a Stratonovich–Weyl (SW) mapping, developed for variable-spin systems [32, 69], in the representation Eq. (2). Each symmetric operator $\hat{O}$ is assigned a Weyl symbol $\mathcal{O}(\bm{\mathcal{Z}})=\mathrm{Tr}[\hat{O}\,\hat{\Delta}(\bm{\mathcal{Z}})]$ and the density matrix is assigned a Wigner function $W(\bm{\mathcal{Z}})=\mathrm{Tr}[\hat{\rho}\,\hat{\Delta}(\bm{\mathcal{Z}})]$ of classical variables $\bm{\mathcal{Z}}$. Expectations are then reproduced as averages, over $W$, of Weyl symbols. The kernel takes the form [68, 33, 8] $\hat{\Delta}(\bm{\mathcal{Z}})=\hat{U}(\Omega)\hat{\Pi}_{\mathcal{J}}\hat{U}^{\dagger}(\Omega)$, for the collective rotation $\hat{U}(\Omega)=e^{-i\hat{J}_{z}\phi}e^{-i\hat{J}_{y}\theta}e^{-i\hat{J}_{z}\psi}$ by Euler angles $\Omega=(\phi,\theta,\psi)$. Here, rotations act on a parity operator $\hat{\Pi}_{\mathcal{J}}$ constructed to describe couplings between elements in Eq. (2) with distinct $J$. The Weyl symbol, ${\mathcal{J}}({\mathcal{J}}+1)$, of $\hat{J}^{2}$ identifies the half-integer $\mathcal{J}=0,1/2,\ldots$ as a classical total spin number analogous to $J$, while the conjugate angle $\psi$ captures mixtures of spin lengths generated by local dissipation, seen by contrasting the Weyl symbols of local and collective jumps, $\hat{l}^{jq}$ and $\hat{L}^{q}$:
@@
Here, $D^{1}_{qj}$ is the spin-1 Wigner-$D$ matrix and $\chi_{jq},\chi_{q}$ are jump amplitudes [61]. For $\gamma^{q}=0$, the spin length is fixed and $\psi$ is redundant, recovering the usual Bloch-sphere description. More generally, $\psi$ is required to capture spin-length mixing under local dissipation. As it is absent in Weyl symbols of collective observables, it serves here as an auxiliary variable in practice [12].
With the quantum–classical mapping specified, we take the semiclassical limit of Eq. (1). For large spin lengths, treating $\mathcal{J}\gg 1$ as continuous yields the semiclassical expansion [32, 69],
@@
up to relative corrections $O(\mathcal{J}^{-1})$, which are typically of order $N^{-1}$. The Poisson bracket $\{\bullet,\bullet\}$ corresponds to a symmetric top on the four-dimensional phase space $\bm{\mathcal{Z}}=(\phi,\theta,\psi,\mathcal{J})$, where the conjugate variable $\psi$ now captures fluctuations in spin length [4]. Substituting Eq. (5) into Eq. (1) yields a truncated-Wigner Fokker–Planck equation for $W$ with positive-semidefinite diffusion [14, 49], which can be equivalently unraveled into Stratonovich stochastic differential (Langevin) equations:
@@
Here $\eta_{k}$ are independent complex white noise processes, $\langle\eta_{k}(t)\eta^{*}_{k^{\prime}}(t^{\prime})\rangle/2=\delta_{kk^{\prime}}\delta(t-t^{\prime})$, with $\circ$ denoting Stratonovich multiplication. $\{\mathcal{L}_{k},r_{k}\}$ are the Weyl symbols and corresponding rates of jump operators, $\mathcal{L}_{k}=\ell^{jq},\mathcal{L}^{q},r_{k}=\gamma^{q},\Gamma^{q}$, for the dissipation channel index $k$.
To finally estimate expectations in the TWA, we average the Weyl symbols over the $N_{\rm traj}$ trajectories $\bm{\mathcal{Z}}^{(i)}(t)$ that sample $W$ at time $t$,
@@
with phase-space variables initialized from the Wigner function of a fully-polarized state [61] and evolved using Eq. (6). In the following, $N_{\mathrm{traj}}=10^{3}\text{--}10^{6}$ typically renders sampling errors of order $1/\sqrt{N_{\mathrm{traj}}}$ negligible, with errors reduced for larger $N$. By evolving only four phase-space variables in parallel per ensemble, this method enables simulation of dynamics in systems from a few emitters to the limit of large ensemble size.
@@
We have numerically verified the accuracy of this general TWA in a number of systems, including boundary time crystals subject to decoherence, and standard superradiant decay [61]. To illustrate the key aspects, we focus here on emission dynamics in the model of a bad-cavity superradiant laser [41, 7], in which collective decay $\Gamma$ (for q=1q=-1) competes with incoherent individual pumping $\gamma$ ($q=1$), and $\hat{H}=0$. These dynamics admit exact Monte Carlo simulations [84], allowing us to verify our results for up to tens of thousands of atoms. While we present two-body observables capturing output emission, we have found that single-body expectations, such as ensemble inversion, are generally predicted with similar accuracy [61].
@@
Fig. 2: Dynamics of an emitter ensemble under local pumping and collective decay. (a) Evolution of normalized emission rate $\langle\hat{J}^{+}\hat{J}^{-}\rangle/N^{2}$ for $N=10,10^{2},10^{3},$ and $10^{6}$ (dark to light), for scaled time $N\Gamma t$. Solid curves show exact results, with TWA shown as dashed curves. For $N=10^{6}$, exact simulation is not feasible here. Inset: maximum error between TWA and exact dynamics for the normalized emission rate (lower line) and normalized inversion $\langle\hat{J}^{z}\rangle/N$ (upper line), compared to a $1/N$ reference (dashed). (b) Squeezing parameter $\xi^{2}$ and subradiance parameter $s=\langle\hat{J}^{+}\hat{J}^{-}\rangle/\langle\hat{J}^{z}+N/2\rangle$ for $N=25$. Inset: steady-state emission rate $\langle\hat{J}^{+}\hat{J}^{-}\rangle/N$ versus pump rate for TWA (dots) and exact solution (solid line); red shaded region indicates values less than one, and the gray region marks the threshold of TWA validity where $\langle\hat{J}^{+}\hat{J}^{-}\rangle$ is of order one.
@@
In Fig. 2(a), we calculate the evolution of normalized intensity $\langle\hat{J}^{+}\hat{J}^{-}\rangle/N^{2}$ for a fully-inverted ensemble as $N$ is increased from ten up to one million. The pump $\gamma=0.25N\Gamma$ is chosen to maintain competition with decay. For ensembles as small as $N=10$, our method accurately captures, with less than 10% deviation, the transient superradiant pulse, subsequent relaxations, and the collective steady-state emission. Crucially, the accuracy increases with $N$. Already for $N\simeq 100$ our predictions agree closely with exact dynamics, and we have quantitatively verified that the prediction error of the TWA decreases with increasing emitter number [Fig. 2(a), inset]. This is consistent with the perturbative expansion Eq. (5), as $\mathcal{J}$ is typically of the order $N$ in individual trajectories. These benchmarks establish the TWA as a reliable semiclassical method for approximating the dynamics of permutation-invariant emitter ensembles.
@@
To demonstrate how the TWA captures nonclassical dynamics, Fig. 2(b) presents the generalized squeezing parameter [70] $\xi^{2}=\langle(\Delta\hat{J}^{x})^{2}+(\Delta\hat{J}^{y})^{2}+(\Delta\hat{J}^{z})^{2}\rangle/(N/2)$ for a weak pump $\gamma=0.4\Gamma$ and moderate $N=25$. A value $\xi^{2}<1$ indicates entanglement relevant for quantum information [40]. The TWA closely reproduces both the transient reduction below unity and the subsequent relaxation to the steady-state. Beyond squeezing, the TWA can also capture subradiance—absent at mean-field level for homogeneous ensembles—where destructive interference suppresses the normalized collective emission rate, $s=\langle\hat{J}^{+}\hat{J}^{-}\rangle/\langle\hat{J}^{z}+N/2\rangle$, below that of independent emitters ($s=1$), offering a distinct resource for metrology [15, 48]. In Fig. 2(b), the TWA accurately predicts the evolution towards $s<1$ and confirms subradiance as a precursor to squeezing in this system [63]. Noticeable deviations from the exact dynamics are confined to the extremely subradiant regime $\langle\hat{J}^{+}\hat{J}^{-}\rangle\lesssim 1$, which coincides with a reduction $\mathcal{J}\sim O(1)$ along individual trajectories—signaling the importance of higher-order terms beyond Eq. (5) and providing a practical diagnostic of accuracy. Away from this regime, the inset of Fig. 2(b) shows that TWA can remain quantitatively accurate across a broad parameter range.
@@
The TWA extends naturally from a single ensemble to composite systems that partition into distinct homogeneous ensembles, including waveguide-coupled emitter arrays, inhomogeneously broadened ensembles, and ensembles in extended resonator lattices [Fig. 3(a)]. By assigning independent phase-space variables $\bm{\mathcal{Z}}_{\ell}$ to each ensemble and evolving Eq. (6), quantum dynamics can be captured with a computational cost that scales linearly with the number of distinct ensembles. As an example, we consider an infinite 1D lattice of coupled resonators in which a subset of $M$ equally spaced sites each hosts a resonant ensemble containing $N$ emitters, for a total $N_{\mathrm{tot}}=NM$. In the bad-cavity limit, adiabatic elimination of the photonic degrees of freedom yields an effective master equation for the collective spins, with resonator-mediated coherent couplings and dissipative jumps:
@@
$\hat{H}=\frac{\Gamma}{2}\sum_{\ell,\ell^{\prime}=1}^{M}\sin\left(\varphi|\ell-\ell^{\prime}|\right)\hat{J}^{+}_{\ell}\hat{J}^{-}_{\ell^{\prime}},\ \ \hat{L}_{\mathrm{F}(\mathrm{B})}=\sum_{\ell=1}^{M}e^{\pm i\varphi\ell}\hat{J}^{-}_{\ell}.$ (8)
@@
In this model, $\Gamma$ denotes the single-emitter decay rate into the chain and $\varphi$ is the phase accumulated between neighboring populated sites. The jump operators $\hat{L}_{\mathrm{F}}$ and $\hat{L}_{\mathrm{B}}$ describe collective emission into forward and backward propagating modes, each at rate $\Gamma/2$. This setup realizes the waveguide-QED paradigm [64] where ensembles act as super-atoms, offering a broad landscape of collective dynamics in the TWA.
@@
Fig. 3: (a) Illustration of composite systems described using the TWA by assigning a collective spin to each permutation-invariant constituent (indicated by color). (b) Forward emission rate $\langle\hat{I}_{\mathrm{F}}\rangle/N_{\mathrm{tot}}^{2}$ for $M=2$ pumped ensembles, each comprising $N=20$ emitters for a total $N_{\mathrm{tot}}=40.$ The propagation phase $\varphi=0,\pi/4,\pi/2$ (upper to lower curves). Emitters are initially in the ground state, and the local pump $\gamma=10\Gamma$. Inset: (identical) single-ensemble emission rates $\langle\hat{I}_{\ell}\rangle/N^{2}$ for $\varphi=0$ as the number of ensembles is increased with $M=1,10,10^{2},10^{3}$ (dark to light) and $\gamma=0.25N_{\mathrm{tot}}\Gamma$. Solid (dashed) curves give the exact solution (TWA).

### v2

@@ Abstract
Emitter ensembles are a fundamental component in quantum optical technologies, yet efficient and accurate simulation of large ensembles remains challenging. Here, we formulate a truncated Wigner approximation (TWA) for permutation-invariant emitter ensembles subject to local dissipation, by sampling stochastic trajectories in an extended phase space encompassing the Bloch sphere. Benchmarks show the TWA can correctly capture dynamics, including nonclassical signatures, with accuracy improving with ensemble size. We demonstrate large-scale simulations of hundreds of interacting ensembles in the TWA to predict emergent spatial coherence and selective directionality of cooperative emission in a pumped 1D chain, highlighting a path to studying extended light-matter systems. Our results expand the scope of scalable simulations of quantum emitter ensembles, establishing a bridge between microscopic models and emergent behavior.
@@
Homogeneous ensembles of emitters play a fundamental role in understanding and engineering open quantum systems. In state-of-the-art experiments featuring atoms [7, 5, 22, 18] and solid-state [63, 90, 40, 55, 28] emitters, collective interactions commonly coexist with local dissipation. While this dissipation must often be mitigated, the interplay with collective dynamics can enable novel dissipative phases of matter [81, 32, 38, 85, 39] and distinct device functionalities enhanced by cooperative behavior [43, 56, 50, 74]. As experiments enter the mesoscopic regime, exact simulations [10, 65, 61, 84, 31, 81, 20] of minimal, permutation-invariant models are typically constrained to modest system sizes. This limitation motivates more scalable approximations [82] such as cumulant expansions [36, 53], tensor networks [78, 92, 87], and (discrete) phase-space methods [62, 45, 68, 25, 23, 44]. However, obtaining quantitatively controlled predictions at larger system sizes remains challenging, calling for targeted approaches [35, 4].
@@
Semiclassical phase-space methods [47, 60] provide a route to approximate many-body quantum dynamics, notably through the truncated Wigner approximation (TWA) [80, 54]. In the TWA, quantum evolution is represented by an ensemble average of semiclassical trajectories. Initial quantum fluctuations enter through random sampling, and in open systems dissipation contributes stochastic evolution. In the semiclassical limit, the approximation improves systematically, with quantum corrections entering perturbatively [54]. While well-established for bosons [70, 17, 79, 6] and for spin ensembles with purely collective dynamics [21, 77, 24, 1], an analogous phase-space approximation of local dissipation in permutation-invariant ensembles has remained limited.
@@
Fig. 1: (Bottom) An ensemble of emitters couples uniformly to a common field. (Top left) The ensemble emits collectively with rate $\Gamma$ and undergoes independent local dissipation with rate $\gamma$. (Top right) In semiclassical simulations, the ensemble spin (large arrow) is sampled by classical phase-space variables (dots). Stochastic trajectories evolve along the Bloch sphere surface under collective processes, while local dissipation drives evolution into the sphere’s interior.
@@
Here, we develop a semiclassical framework for the dynamics of permutation-invariant emitter ensembles (Fig. 1). Building on a quantum-to-classical mapping for variable-length spins [33], we formulate a truncated Wigner approximation in which the entire ensemble is described using four phase-space variables that extend the usual Bloch sphere description. The additional variables capture changes in total spin induced by local dissipation, which enters through a few stochastic processes acting on semiclassical trajectories. In bad-cavity superradiance, and boundary time crystals, we find the method performs well for small ensembles of tens of emitters—capturing squeezing and subradiance—and numerically approaches the exact solution for increasing ensemble size. The phase-space dimension scales linearly with the number of distinct ensembles, enabling simulations corresponding to tens of thousands of emitters interacting through a resonator chain. In this extended nonlinear medium, we illustrate spatially structured dynamics under uniform pumping and a crossover to strong, selectively directional emission in response to a weak probe, highlighting emergent cooperative phenomena in light-matter systems accessible within the TWA.
@@
We consider a homogeneous ensemble formed from $N$ two-level emitters indexed by $n$. Coupling to common modes induces coherent evolution under a Hamiltonian $\hat{H}$ and collective dissipation via jumps $\hat{L}^{q}$ built from collective spin operators $\hat{L}^{0}=\hat{J}^{z}=\frac{1}{2}\sum_{n}\hat{\sigma}_{n}^{z},\ \hat{L}^{\pm 1}=\hat{J}^{\pm}=\sum_{n}\hat{\sigma}_{n}^{\pm}$ where $\hat{\sigma}^{0}_{n}\equiv\hat{\sigma}^{z}_{n}$, $\hat{\sigma}^{\pm 1}_{n}\equiv\hat{\sigma}^{\pm}_{n}$ are the Pauli-$z$ and ladder operators of emitter $n$. Including statistically identical, independent dissipation, the system evolves under the Born–Markov master equation ($\hbar=1$),
@@
To obtain a semiclassical approximation of the dynamics in Eq. (1), we use the equivalent angular momentum representation [65, 3, 84, 64]. Here, any permutation-invariant density matrix can be expressed,
@@
In this representation, $|JM\rangle$ denotes collective kets labeled by eigenvalues $J(J+1)$ and $M$ of total spin squared $\hat{J}^{2}$ and inversion $\hat{J}^{z}$, respectively. Elements $|JM\rangle\langle JM^{\prime}|$ correspond to degeneracy-averaged Dicke-state outer products in the full Hilbert space. $\hat{\rho}$ is then block diagonal in $J$, and collective operators $\hat{J}^{q}$ act as rotations in each block. Local dissipation induces population transfer between distinct sectors $J\leq N/2$ and, crucially, can be expressed [35] using effective jumps $\hat{l}^{jq}$ in the representation Eq. (2):
@@
The three jump operators $\hat{l}^{jq}$ (j=1,0,1j=-1,0,1) for each $q$, couple both inversion $M\to M+q$ and total spin number $J\to J+j$ acting on elements $|JM\rangle$, completing the angular momentum representation of Eq. (1).
In the representation Eq. (2), we establish a quantum–classical mapping (details in [64]) for operators in Eq. (1), by applying a Stratonovich–Weyl (SW) correspondence developed for variable-spin systems [33]. For the SW kernel $\hat{\Delta}$, each operator $\hat{O}$ is assigned a Weyl symbol $\mathcal{O}(\bm{\mathcal{Z}})=\mathrm{Tr}[\hat{O}\,\hat{\Delta}(\bm{\mathcal{Z}})]$ and the density matrix is assigned a Wigner function $W(\bm{\mathcal{Z}})=\mathrm{Tr}[\hat{\rho}\,\hat{\Delta}(\bm{\mathcal{Z}})]$ of classical variables $\bm{\mathcal{Z}}$. Expectations are then reproduced as averages over $W$ of Weyl symbols. For a collective spin, the kernel takes the covariant form [71, 34, 8] $\hat{\Delta}(\bm{\mathcal{Z}})=\hat{U}(\bm{\Omega})\hat{\Pi}_{\mathcal{J}}\hat{U}^{\dagger}(\bm{\Omega})$, for the collective rotation $\hat{U}(\bm{\Omega})$ by Euler angles $\bm{\Omega}=(\phi,\theta,\psi)$. Here, $\hat{U}$ acts on a generalized parity operator $\hat{\Pi}_{\mathcal{J}}$ that incorporates couplings between elements in Eq. (2) with distinct $J$. The Weyl symbol, ${\mathcal{J}}({\mathcal{J}}+1)$, of $\hat{J}^{2}$ identifies the additional half-integer index $\mathcal{J}=0,1/2,\ldots$ as a classical total spin number analogous to $J$. To elucidate the role of $\psi$, we calculate Weyl symbols of local (collective) operators $\hat{l}^{jq}$ ($\hat{L}^{q}$):
@@
Here, $D^{1}_{qj}$ is the spin-1 Wigner-$D$ matrix and $\chi$ represents jump amplitudes. Crucially, dependence on $\psi$ arises only in matrix elements with $j\neq 0$, corresponding to jumps $\hat{l}^{jq}$ that change total spin. While spherical angles $(\phi,\theta)$ on the Bloch sphere capture strictly collective dynamics for $\gamma^{q}=0$, variables $(\psi,\mathcal{J})$ are then required to describe mixing of total spin under local dissipation. As $\psi$ is absent from Weyl symbols of collective observables, it serves as an auxiliary variable in practice [12].
With a complete quantum–classical mapping specified in the representation Eq. (2), we apply a semiclassical expansion of Eq. (1). Treating $\mathcal{J}\gg 1$ as continuous yields the approximation [72],
@@
where the Poisson bracket $\{\bullet,\bullet\}$ corresponds to a linear rigid rotor for the four-dimensional phase-space coordinates $\bm{\mathcal{Z}}=(\phi,\theta,\psi,\mathcal{J})$ (see Appendix). When the total spin is macroscopic, relative corrections $O(\mathcal{J}^{-1})$ are typically of order $N^{-1}$, providing a semiclassical control parameter. Applying the SW correspondence to Eq. (1) and then using Eq. (5) yields a Fokker–Planck equation for $W$ with nonnegative diffusion [14, 64], which we unravel into a system of stochastic differential equations:
@@
Here, $\eta_{k}$ are independent complex white noise processes, $\langle\eta_{k}(t)\eta^{*}_{k^{\prime}}(t^{\prime})\rangle/2=\delta_{kk^{\prime}}\delta(t-t^{\prime})$, with $\circ$ denoting Stratonovich multiplication. $(\mathcal{L}_{k},r_{k})$ are the Weyl symbols and corresponding rates of jump operators, $\mathcal{L}_{k}=\ell^{jq},\mathcal{L}^{q},r_{k}=\gamma^{q},\Gamma^{q}$, for the dissipation channel index $k$, while $\mathcal{H}$ is the Weyl symbol of the Hamiltonian.
Finally, to estimate observable expectations for the dynamics in Eq. (1) using the semiclassical evolution Eq. (6), we average Weyl symbols over $N_{\mathrm{traj}}$ trajectories $\bm{\mathcal{Z}}^{(i)}(t)$ that approximately sample $W$ at time $t$,
@@
with phase-space variables initialized by approximating the Wigner function of a fully-polarized state [64] and evolved using Eq. (6). In the following, $N_{\mathrm{traj}}=10^{3}\text{--}10^{6}$ typically renders sampling errors of order $1/\sqrt{N_{\mathrm{traj}}}$ negligible, with errors reduced for larger $N$. By evolving four phase-space variables in parallel per ensemble, this TWA enables semiclassical simulation of dynamics in systems from a few emitters to the limit of large ensemble size.
@@
We have numerically verified the accuracy of this general TWA in a number of systems, including boundary time crystals subject to decoherence (see Appendix) and standard superradiant decay [64]. To illustrate the key aspects, we focus here on emission dynamics in the model of a bad-cavity superradiant laser [43, 7], in which collective decay $\Gamma$ (for q=1q=-1) competes with incoherent individual pumping $\gamma$ ($q=1$), and $\hat{H}=0$. These dynamics admit exact Monte Carlo simulations [88], allowing us to verify the TWA for thousands of emitters in the following. While we present two-body observables capturing emission, we have found that single-body expectations, such as ensemble inversion, are generally predicted with similar accuracy.
@@
Fig. 2: Dynamics of a permutation-invariant ensemble under local pumping and collective decay. (a) Evolution of normalized emission rate $\langle\hat{J}^{+}\hat{J}^{-}\rangle/N^{2}$ for $N=10,10^{2},10^{3},$ and $10^{6}$ (dark to light), for scaled time $N\Gamma t$ (axis break as indicated). Solid curves show exact results, with TWA shown as dashed curves. For $N=10^{6}$, exact simulation is not feasible here. Inset: maximum absolute error between TWA and exact dynamics for the normalized emission rate (lower line) and normalized inversion $\langle\hat{J}^{z}\rangle/N$ (upper line) over the full evolution, compared to a $1/N$ reference (dashed). (b) Squeezing parameter $\xi^{2}$ and subradiance parameter $s$ for $N=25,\gamma=0.4\Gamma$. Inset: steady-state emission rate $\langle\hat{J}^{+}\hat{J}^{-}\rangle/N$ versus pump rate for TWA (dots) and exact solution (solid line); the red shaded region indicates values less than 1, and gray marks the threshold of TWA validity where $\langle\hat{J}^{+}\hat{J}^{-}\rangle$ is of order 1.
@@
In Fig. 2(a), we calculate the evolution of normalized emission rate $\langle\hat{J}^{+}\hat{J}^{-}\rangle/N^{2}$ for a fully-inverted ensemble as $N$ is increased from ten up to one million. The pump $\gamma=0.25N\Gamma$ is chosen to maintain competition with decay. For ensembles as small as $N=10$, our method captures, with less than 10% deviation, the transient superradiant pulse, subsequent relaxations, and the collective steady-state emission. Crucially, the accuracy increases with $N$. Already for $N\simeq 100$, our predictions agree closely with exact dynamics, and we have quantitatively verified that the prediction error of the TWA decreases with increasing emitter number [Fig. 2(a), inset]. This is consistent with the perturbative expansion Eq. (5), as $\mathcal{J}$ is of order $N$ in individual trajectories. These benchmarks establish the TWA as a semiclassical method for approximating the dynamics of permutation-invariant emitter ensembles.
@@
To demonstrate how the TWA captures nonclassical dynamics, Fig. 2(b) presents the generalized squeezing parameter [73] $\xi^{2}=\langle(\Delta\hat{J}^{x})^{2}+(\Delta\hat{J}^{y})^{2}+(\Delta\hat{J}^{z})^{2}\rangle/(N/2)$ for a weak pump $\gamma=0.4\Gamma$ and moderate $N=25$. A value $\xi^{2}<1$ indicates entanglement relevant for quantum information [42]. The TWA closely reproduces both the transient reduction below unity and the subsequent relaxation to the steady-state. In addition to squeezing, the TWA can also capture subradiance—absent in a mean-field description of permutation-invariant ensembles. Here, destructive interference suppresses the normalized collective emission rate, $s=\langle\hat{J}^{+}\hat{J}^{-}\rangle/\langle\hat{J}^{z}+N/2\rangle$, below that of independent emitters ($s=1$), offering a distinct resource for metrology [15, 50]. In Fig. 2(b), the TWA closely tracks the evolution towards $s<1$ and confirms subradiance as a precursor to squeezing in this system [66]. Noticeable deviations from the exact dynamics are confined to the extremely subradiant regime $\langle\hat{J}^{+}\hat{J}^{-}\rangle\lesssim 1$, which coincides with a reduction $\mathcal{J}\sim O(1)$ along individual trajectories—signaling the importance of higher-order terms beyond Eq. (5) and providing a practical diagnostic of accuracy. Away from this regime, the inset of Fig. 2(b) shows that TWA can remain quantitatively accurate across a broad parameter range.
@@
The TWA extends naturally from a single ensemble to composite systems that partition into distinct permutation-invariant ensembles, including waveguide-coupled emitter arrays, inhomogeneously broadened ensembles, and ensembles in extended resonator lattices [Fig. 3(a)]. By assigning independent phase-space variables $\bm{\mathcal{Z}}_{m}$ to each ensemble and evolving using Eq. (6), quantum dynamics can be captured by a set of phase-space equations whose dimensionality scales linearly with the number of distinct ensembles (see Appendix). As an example, we consider an infinite 1D lattice of coupled resonators in which $M$ equally spaced sites each host a resonant ensemble containing $N$ emitters, for a total $N_{\mathrm{tot}}=NM$. In the bad-cavity limit, adiabatic elimination of the photonic degrees of freedom yields an effective master equation for the emitter ensembles, with resonator-mediated coherent couplings and dissipative jumps:
@@
$\hat{H}=\frac{\Gamma}{2}\!\sum_{\begin{subarray}{c}m=1\\ m^{\prime}=1\end{subarray}}^{M}\!\sin\left(\varphi|m{-}m^{\prime}|\right)\hat{J}^{+}_{m}\hat{J}^{-}_{m^{\prime}},\ \hat{L}_{\mathrm{F}(\mathrm{B})}\!=\!\sum_{m=1}^{M}e^{\pm i\varphi m}\hat{J}^{-}_{m}.$ (8)
@@
In this model, $\Gamma$ denotes the decay rate into the chain and $\varphi$ is the phase accumulated between neighboring populated sites. The jump operators $\hat{L}_{\mathrm{F}}$ and $\hat{L}_{\mathrm{B}}$ describe collective emission into forward (++) and backward (-) propagating modes, each at rate $\Gamma/2$. This setup realizes the waveguide-QED paradigm [67] where ensembles act as super-atoms, offering a broad landscape of collective dynamics in the TWA.
@@
Fig. 3: (a) Illustration of composite systems described using the TWA by assigning a distinct collective spin (arrow) to each permutation-invariant emitter ensemble (dots). (b) Normalized forward emission rate $\langle\hat{I}_{\mathrm{F}}\rangle/N_{\mathrm{tot}}^{2}$ for $M=2$ pumped ensembles, each comprising $N=20$ emitters for a total $N_{\mathrm{tot}}=40,$ in the system Eq. (8). The propagation phase $\varphi=0,\pi/4,\pi/2$ (upper to lower curves). Emitters are initially in the ground state, and the local pump $\gamma=10\Gamma$ for each ensemble. Inset: (identical) site-resolved emission rates $\langle\hat{I}_{m}\rangle/N^{2}$ for $\varphi=0$ as the number of ensembles is increased with $M=1,10,10^{2},10^{3}$ (dark to light) and $\gamma=0.25N_{\mathrm{tot}}\Gamma$. Solid (dashed) curves give the exact solution (TWA).